# State Estimation Using Sparse DEIM and Recurrent Neural Networks

### v2

@@ 4 Numerical Results
Table 1: Parameters for each test case.
@@
$\mbox{RE}=\frac{|\widetilde{\mathbf{u}}-\mathbf{u}|}{|\mathbf{u}-\overline{\mathbf{u}}|},$ (22)
where $\mathbf{u}$ is the true state, $\widetilde{\mathbf{u}}$ is the estimation, and $\overline{\mathbf{u}}$ is the empirical mean of the data set from which the POD modes are computed. Note that in the numerator, the mean cancels out and therefore it is more prudent to remove the mean from the denominator as well, in order to avoid misleadingly small relative errors.
@@ 4.1 Lorenz-96 model
Figure 5(a) shows the relative error over the time interval $t\in[0,200]$. The Q-DEIM error oscillates around 77%, indicating a poor estimation. In contrast, the S-DEIM error is around 29% which is close to that of the best possible fit at around 21%. We recall that the RC network takes some time to synchronize to the observational input data. This transient phase is evident in figure 5(b) which shows a close-up view of the relative error. After about two time units, equivalent to $8\Delta t$, the synchronization is complete and the S-DEIM relative error has settled around $29\%$.
@@ 4.2 Kuramoto–Sivashinsky equation
where the solution $u(x,t)$ is defined on the spatial domain $x\in[-L/2,L/2]$ with periodic boundary conditions. A larger domain size $L$ typically leads to more complicated spatiotemporal dynamics. Here, we set $L=22$ which is known to lead to chaotic dynamics [10]. We use a standard pseudo-spectral method with $N=128$ collocation points to solve the KS equation.
@@ 4.3 Rayleigh–Bénard convection
As our last example, we demonstrate the application of S-DEIM on the two-dimensional Rayleigh–Bénard convection. This problem describes the motion of a fluid inside a rectangular domain $[0,L]\times[0,H]$. As depicted in figure 8, the bottom wall is kept at a relatively high temperature $T_{b}$ and the top wall is maintained at a lower temperature $T_{t}$, where $T_{b}>T_{t}$. The fluid is heated by the bottom wall and rises up towards the top wall. The fluid is subsequently cooled down and descends towards the bottom wall again. This process leads to convection cells which can exhibit chaotic dynamics [34, 27].
@@
where $T(x,z,t)$ denotes the fluid temperature and $\mathbf{v}(x,z,t)$ is the fluid velocity field. This model has two dimensionless parameters: Rayleigh number $Ra$ and Prandtl number $Pr$. Throughout this section, we set $Ra=10^{5}$ and $Pr=10$. For the domain size, we use $L=4$ and $H=1$. The relevant boundary conditions are listed in figure 8. In our numerical simulations, we set $T_{b}=1$ and $T_{t}=0$ for the wall temperatures.
@@
We seek to estimate the temperature field $T(x,z,t)$ from its sparse observations. Figure 9 shows the relative errors of state estimations using $r=25$ sensors and $m=50$ POD modes. The Q-DEIM estimates, which use the trivial kernel vector $\mathbf{z}=\mathbf{0}$, have a mean relative error of approximately $68\%$. The S-DEIM reconstruction, which uses an RNN to estimate the optimal kernel vector, achieves much smaller errors. Recall that RNNs require some time to synchronize with the input data. As such, the S-DEIM errors are initially large. But within approximately $15$ time units, equivalent to $30\Delta t$, the error decays towards its mean relative error of approximately $26\%$.
@@ 5 Conclusions
We proposed an equation-free method for estimating the state of a spatiotemporal dynamical system from incomplete observations. Our method relies on Sparse Discrete Empirical Interpolation Method (S-DEIM) which comprises two terms. The first term can immediately be evaluated using the observations. The second term involve the kernel vector of a linear operator which cannot be estimated from instantaneous observational data. Here, we developed a machine learning method to estimate this optimal kernel vector. More precisely, we train a recurrent neural network whose inputs are the time series of the observations (including past history) and its output estimates the optimal kernel vector of S-DEIM.
We demonstrated on three numerical examples that S-DEIM, together with the RNN training, returns nearly optimal state estimations. Furthermore, the computational cost of the proposed method is exceedingly low. In each numerical example, the training process took only a few seconds. Using the trained networks to obtain a state estimation took only a fraction of a second.
@@
Acknowledgments.
I am grateful to Zack Hilliard for help with setting up the numerical simulations for the Rayleigh–Bénard convection. This work was supported by the National Science Foundation, the Algorithms for Threat Detection (ATD) program, through the award DMS-2220548.

### v3

@@ 4 Numerical Results
Table 2: Parameters for each test case. The bottom half of the table reports the information regarding the RC network.
@@
Reservoir size ($K$) 200 300 500
Reservoir sparsity 10% 10% 10%
Training/Inference time (seconds) 0.2/0.02 0.7/0.04 0.3/0.02
@@
$\mbox{RE}=\frac{\|\widetilde{\mathbf{u}}-\mathbf{u}\|}{\|\mathbf{u}-\overline{\mathbf{u}}\|},$ (23)
where $\mathbf{u}$ is the true state, $\widetilde{\mathbf{u}}$ is the estimation, and $\overline{\mathbf{u}}$ is the empirical mean of the data set from which the POD modes are computed. Note that in the numerator, the mean $\overline{\mathbf{u}}$ cancels out and therefore it is more prudent to subtract the mean from the denominator as well, as in equation (23). Otherwise, a large mean will result in misleadingly small relative errors.
@@ 4.1 Lorenz-96 model
Figure 5(a) shows the relative error over the time interval $t\in[0,200]$. The Q-DEIM error oscillates around 77%, indicating a poor estimation. In contrast, the S-DEIM error is around 29% which is close to that of the best possible fit at around 21%. It is remarkable that although both methods use the same observational data, S-DEIM reduces the reconstruction error by 48% compared to Q-DEIM.
We recall that the RC network takes some time to synchronize to the observational input data. This transient phase is evident in figure 5(b) which shows a close-up view of the relative error. After about two time units, equivalent to $8\Delta t$, the synchronization is complete and the S-DEIM relative error has settled around $29\%$.
@@ 4.2 Kuramoto–Sivashinsky equation
where the solution $u(x,t)$ is defined on the spatial domain $x\in[-L/2,L/2]$ with periodic boundary conditions. A larger domain size $L$ typically leads to more complicated spatiotemporal dynamics. Here, we set $L=22$ which is known to lead to chaotic dynamics [46]. We use a standard pseudo-spectral method with $N=128$ collocation points to solve the KS equation [47].
@@ 4.3 Rayleigh–Bénard convection
As our last example, we demonstrate the application of S-DEIM on the two-dimensional Rayleigh–Bénard convection. This problem describes the motion of a fluid inside a rectangular domain $0,L]\times[0,H]$. As depicted in figure 8, the bottom wall is kept at a relatively high temperature $T_{b}$ and the top wall is maintained at a lower temperature $T_{t}$, where $T_{b}>T_{t}$. The fluid is heated by the bottom wall and rises up towards the top wall. The fluid is subsequently cooled down and descends towards the bottom wall again. This process leads to convection cells which can exhibit chaotic dynamics [48, 49].
@@
where $T(x,z,t)$ denotes the fluid temperature, $p(x,z,t)$ denotes pressure, and $\mathbf{v}(x,z,t)$ is the fluid velocity field. The unit vector $\mathbf{e}_{z}$ points in the opposite direction of gravity. This model has two dimensionless parameters: Rayleigh number $Ra$ and Prandtl number $Pr$. Throughout this section, we set $Ra=10^{5}$ and $Pr=10$. For the domain size, we use $L=4$ and $H=1$. The relevant boundary conditions are listed in figure 8. In our numerical simulations, we set $T_{b}=1$ and $T_{t}=0$ for the wall temperatures. A Fourier pseudo-spectral method is used for discretizing the PDEs in space and a 4th order Runge-Kutta scheme is used for temporal discretization [50, 51].
@@
We seek to estimate the temperature field $T(x,z,t)$ from its sparse observations. The velocity field $\mathbf{v}(x,z,t)$ is not observed or used in these estimations. Figure 9 shows the relative errors of state estimations using $r=25$ sensors and $m=50$ POD modes. The Q-DEIM estimates, which use the trivial kernel vector $\mathbf{z}=\mathbf{0}$, have a mean relative error of approximately $68\%$. The S-DEIM reconstruction, which uses an RNN to estimate the optimal kernel vector, achieves much smaller errors. Recall that RNNs require some time to synchronize with the input data. As such, the S-DEIM errors are initially large. But within approximately $15$ time units, equivalent to $30\Delta t$, the error decays towards its mean relative error of approximately $26\%$.
@@
4.4 Comparison with direct learning
Recall the best fit approximation (4) which corresponds to the expansion coefficients $\hat{\boldsymbol{\alpha}}(t)=\Phi_{m}^{\top}\mathbf{u}(t)$. Although these coefficients are not directly computable from observations, we can nonetheless try to estimate them from observational time series $\mathcal{Y}(t)$ using an RNN. More precisely, we can train an RNN whose inputs are the observational time series $\mathcal{Y}(t)$ (similar to S-DEIM) and its outputs are the best fit coefficients $\hat{\boldsymbol{\alpha}}(t)$ (unlike S-DEIM where the outputs are the optimal kernel vector $\hat{\boldsymbol{\xi}}(t)$). We refer to this approach as the direct learning method and in this section compare its performance to S-DEIM.
The direct method has two drawbacks when compared to S-DEIM. First, the dimension of the RNN output space is larger for the direct method. For the direct method, the outputs $\hat{\boldsymbol{\alpha}}$ are $m$-dimensional, whereas in S-DEIM the outputs $\hat{\boldsymbol{\xi}}$ are (mr)(m-r)-dimensional. This may render the training of the RNN for the direct method slightly more challenging. But, the more important issue is the loss of the interpolation property (17). Owing to the interpolation property, S-DEIM estimates agree exactly with the observations. In contrast, there is no such guarantee for the direct method.
In spite of these drawbacks, the direct method should be theoretically more accurate since it seeks to estimate the best fit. However, as we show in figure 11 this higher accuracy does not materialize in practice. The top row of figure 11 shows the total relative error (23) for the direct learning method and S-DEIM. The total relative errors for both methods are remarkably similar, across all three examples: Lorenz-96, KS, and RBC systems. Here, to perform a fair comparison, we used the same RC architecture for the direct learning method and for S-DEIM, except for the output dimensions. Furthermore, both methods are trained on the same input data and also tested on identical datasets. Figure 11 shows the results on the test data.
Figure 11: Comparison of S-DEIM and direct learning. The top row shows the total relative error (23). The bottom row shows the relative observation error. The numbers report the time-average of the observation errors for the direct method.
As mentioned earlier, S-DEIM is guaranteed to satisfy the interpolation property. This is shown in the bottom row of figure 11 where the relative observation error $\|S_{r}^{\top}\widetilde{\mathbf{u}}-\mathbf{y}\|/\|\mathbf{y}\|$ is plotted. The S-DEIM observation error is about $10^{-13}\%$ for all three systems; it is not exactly zero because of numerical round-off errors. The observation error for the direct method is at least thirteen orders of magnitude larger, ranging between 6.8-30% on average.
In summary, we observe that the direct method not only loses the interpolation property but also fails to improve the estimation error appreciably. As a result, we do not recommend using the direct learning method instead of S-DEIM.
@@ 5 Conclusions
We proposed an model-free method for estimating the state of a spatiotemporal dynamical system from incomplete observations. Our method relies on Sparse Discrete Empirical Interpolation Method (S-DEIM) which comprises two terms. The first term can immediately be evaluated using the observations. The second term involve the kernel vector of a linear operator which cannot be estimated from instantaneous observational data. Here, we developed a machine learning method to estimate this optimal kernel vector. More precisely, we train a recurrent neural network whose inputs are the time series of the observations (including past history) and its output estimates the optimal kernel vector of S-DEIM.
We demonstrated on three numerical examples that S-DEIM, together with the RNN training, returns state estimations with nearly optimal kernel vectors. The resulting S-DEIM estimations are 42-58% more accurate than the Q-DEIM estimates. Furthermore, the computational cost of the proposed method is exceedingly low. In each numerical example, the training process took only a few seconds. Using the trained networks to obtain a state estimation took only a fraction of a second.
With certain adjustments, our method is scalable to large scale systems, such as three-dimensional turbulent flows, where the state space dimension $N$ is extremely large. The bulk of the computational cost is associated with the offline computations, as depicted in figure 2. Within this stage, producing the high-fidelity experimental or simulation data and the subsequent POD computation are the most resource intensive steps. For instance, in large scale applications, it is computationally infeasible to compute POD modes from a direct singular value decomposition. Instead, one should use low cost and low memory methods such as randomized matrix decomposition [52] or tensor-based methods [21]. It is worth mentioning that the training phase is not significantly impacted by the size $N$ of the state space. Even in large scale systems, the input and output dimensions of the RNN are determined by the number of sensors $r$ and the number of modes $m$, which are relatively small.
@@
Data and code
The codes for this work can be found at https://github.com/mfarazmand/S-DEIM.git and the corresponding data is available at
https://doi.org/10.5281/zenodo.15832022.
Acknowledgments
This work was supported by the National Science Foundation, the Algorithms for Threat Detection (ATD) program, through the award DMS-2220548. I am grateful to Zack Hilliard for help with setting up the numerical simulations for the Rayleigh–Bénard convection. I would also like to thank Prof. Arvind Saibaba for his comments on an earlier version of this paper.